# Generating Fine Details of Entity Interactions

### v1

@@ 5 Experiments
We compare DetailScribe with state-of-the-art text-to-image generation models on the InterActing dataset. Additionally, we carry out ablation studies on the number of re-denoising refinement steps and the incorporation of hierarchical concept decomposition.
Figure 4: Images generated by DetailScribe and baselines on the InterActing dataset. From left to right: 1) Stable Diffusion: generating images using Stable Diffusion (SD) with the prompt directly; 2) SD + GPT: Stable Diffusion with GPT augmented prompts ; 3) DALL·E 3: prompting DALL·E 3 with the original prompts, which are augmented internally within DALL·E; 4) Ours: DetailScribe generating images with decomposed concepts and VLM generated critiques. DetailScribe consistently provides effective corrections, which help generate images that closely follow the fine details in the prompts.
@@ 5.1 Baselines and Implementation Details
Functional Relation Compositional Relation Multi-subject Interaction
Human GPT-4o ImReward CLIPS. B-VQA Human GPT-4o ImReward CLIPS. B-VQA Human GPT-4o ImReward CLIPS. B-VQA
SD 3.360 4.680 1.657 0.971 0.366 3.583 4.533 1.415 0.898 0.379 3.225 4.000 1.171 0.894 0.306
+ GPT Rewrite 3.770 4.680 1.546 0.921 0.245 3.167 4.533 1.283 0.867 0.325 3.275 4.200 1.101 0.890 0.292
+ GPT Refine 3.450 4.200 1.524 0.951 0.290 3.667 4.867 1.390 0.889 0.389 3.175 3.700 1.022 0.858 0.292
+ Multi-seed 3.270 4.560 1.718 0.985 0.434 3.650 4.733 1.538 0.912 0.423 3.375 4.000 1.149 0.903 0.302
DALL·E 3 3.940 4.720 1.535 0.880 0.226 3.433 4.867 1.382 0.838 0.367 3.775 4.600 1.111 0.813 0.286
DetailScribe 4.280 4.960 1.761 0.998 0.449 4.283 5.000 1.545 0.923 0.485 3.800 4.400 1.326 0.907 0.343
Table 2: Average human/VLM likert scale (1 - 5) and pre-trained metrics on three scenarios of sampled InterActing dataset. We report the human Likert scale (Human Evaluation), VLM evaluation score (GPT-4o), as well as ImageReward (ImReward), CLIPScore (CLIPS.) and BLIP-VQA (B-VQA) score. DetailScribe receives the highest scores according to human preference in all scenarios.
We compared our approach with the state-of-the-art text-to-image generation models as follows:
  • Stable Diffusion. We generate the image conditioned on the prompts in InterActing using the Stable Diffusion SD3.5-large model (SD). We used the SD3.5-large model [4] for all the baselines that are based on a pre-trained T2I generative model.
  • Refinement-Augmented Generation We also include two common strategies for refinement-augmented generation (SD + GPT Rewrite and SD + GPT Refine). For SD + GPT Rewrite, we first use GPT-4o to generate a detailed text prompt given the initial prompt from InterActing, akin to the concept decomposition used in DetailScribe. This improved text prompt is then used to re-generate images with the same pre-trained T2I model. For SD + GPT Refine, we follow the strategy in [27]. Specifically, GPT-4o was adopted to provide a refined prompt based on an initially generated image and the refinement request prompt which is used in [27]. We then re-generated the image with the refined prompt.
  • Inference scaling [21]. We implemented a toy version of noise searching in [21] (SD + Multi-seed). Specifically, we sampled two different noise for image generation. We adopted CLIPScore [7] as its verifier.
  • DALL·E [24, 1] internally integrates LLMs (GPT) to refine prompts with detail model interprets effectively before generating images. We include the images generated by DALL·E 3 as a strong baseline to assess DetailScribe’s advancement in interpreting and generating scenes with rich entity interactions.
  • DetailScribe. Our DetailScribe implementation leverages Stable Diffusion 3.5 as the foundational model for both image generation and refinement. To ensure fair comparisons, all approaches involving VLMs and LLMs use separate, identical prompts with the same GPT-4o model, maintaining consistency across evaluations.
Among all the algorithms, SD + GPT Refine and inference scaling (SD + Multi-seed), as well as DetailScribe, require two times the computation of the base model (SD) for one iteration of refinement (We neglected the critique time of the VLM done by commercial APIs, which has a runtime of 10% of SD3.5.) SD + GPT Rewrite refines prompt unconditional on previous generation, thus requires the same computation as SD.
@@ 5.2 Result
We evaluate the models on three scenarios from the InterActing dataset and report the results separately. Due to the scalability of high-quality human evaluation, we sampled 50 prompts from InterActing for both human evaluation and automatic evaluation (Table 2), and compared the agreement in between. Overall, VLM evaluator achieves the highest agreement at 90.4%, compared to the other metrics: ImageReward (73.6%), CLIPScore (70.4%), and BLIP-VQA (67.6%). We further presented the automatic evaluation on the entire InterActing in Table 3. DetailScribe outperformed all methods based on SD3.5 in all evaluation.
GPT-4o ImReward CLIPS. B-VQA
SD 4.107 1.323 0.902 0.336
+ GPT Rewrite 4.021 1.193 0.880 0.268
+ GPT Refine 3.999 1.255 0.880 0.300
+ Multi-seed 4.126 1.354 0.922 0.365
DALL·E 3 4.496 1.222 0.860 0.312
DetailScribe 4.557 1.460 0.923 0.381
Table 3: Automatic evaluation on entire InterActing. DetailScribe outperforms all baselines on all pre-trained metrics. We include evaluation by scenario in Appendix B.
Fig. 4 shows more examples generated by our model and the baselines. As shown in the figure, DetailScribe is able to generate images with fine details delineating entity interactions. For example, the second row demonstrates the capability of DetailScribe in capturing functional relations. Given the prompt “A cat sails across the sea in a large seashell, holding a mast.”, all baselines fail to capture the relation “holding”, while DetailScribe is able to generate accurate details of a cat holding the mast with fine details. Similarly, as shown in the first row, DetailScribe is able to depict the “rolling” relation accurately through the critique-and-refinement process, while the SD3.5 model with original prompt and GPT refined prompt place the hand of the hedgehog at an unrealistic location on the rolling pin. The 5th row contains a challenging example of a complex scene layout “zig-zag path”. DetailScribe is the only model capable of generating such fine layout patterns. Stable Diffusion can generate an image with zigzag patterns but fails to reveal a path. Both SD with GPT rewritten prompt and DALL·E fail to follow the prompt on the zigzag pattern and only generate a path with leaves.
@@ 5.3 Ablation: Hierarchical Concept Decomposition Improves Error Detection
Figure 5: An illustrative example showing the effectiveness of the explicit concept decomposition module. VLM first critiques the original generation, and identifies the features needs to be correct (red) and the features non-necessary for further modification (green), and then provides the corrected prompt for re-denoising.
Human GPT-4o ImReward CLIPS. B-VQA
w/o. Decomp 3.843 4.720 1.586 0.953 0.410
w/. Decomp 4.187 4.840 1.609 0.957 0.438
Table 4: Ablation study on the effectiveness of the concept decomposition module. Including explicit concept decomposition significantly improves the generation quality.
@@
Fig. 5 shows an illustrative example. The top box in the figure shows critiques generated without our decomposed concepts. The VLM with access to only the initial prompt does not properly attend to the detailed properties. Instead, the critiques focus more on global attributes such as the shape of objects, lighting conditions, or object arrangements. With the concept decomposition step explicitly added, the VLM can generate better critiques by attending to local details such as “missing a spoon” or action concepts such as “stirring”. With more errors detected and included in the refined prompt, we also see an improvement in the re-denoising image refinement.
@@ 5.4 Ablation: Re-Denoising Step
We also study the impact of using different numbers of re-denoising steps, as shown in Fig. 6. We generate images with SD3.5, starting from steps $T$, $T-1$, $T-4$, and $T-6$, respectively. Introducing the refinement prompt at a later de-noising step results in images that are more similar to the original ones, as the diffusion model has fewer steps to refine the generated results. Overall, we find that salient and global attributes of objects, such as the shapes and colors of large objects, are less likely to be modified if the re-denoising occurs at a late stage. However, it is still possible to make local changes that do not interact with large regions of the image, such as adding small objects.
We also evaluate the performance of generating images from pure noise using the refined prompt. Given that the refined prompts contain more details, we empirically observe that diffusion models are more prone to missing concepts or concept leakage, such as missing entities (e.g. ignoring “Snow Bunny” given the prompt “A Rabbit Sculpting Snow Bunny”). Thus, our refinement-augmented generation procedure can also be interpreted as a coarse-to-fine generation process, making it easier for the model to generate images with coherent global structures and detailed local attributes simultaneously. We provide more examples in the supplementary material. Furthermore, while we find that a one-round refinement is sufficient, future work may explore extending the framework to multi-round refinements.
Figure 6: Ablation on the number of re-denoising steps. The first column shows the result without the refined prompt. Column 2-5: Starting the re-denoising at step $T$, $T-1$, $T-4$, and $T-6$.
Due to the same reason that prevents the re-denoising step from introducing large-patch changes to the image content, a current limitation of the DetailScribe framework is its assumption that the image generated without prompt refinement has a correct global scene structure. For example, if the generated image based on the user input completely misses one of the main subjects. In such cases, even if the VLM detects errors in the generated image, the re-denoising process is not capable of fixing them. Future work may explore seed search [21] based on similar critiques strategy.
@@ 6 Conclusion
In this paper, we introduce InterActing, a comprehensive dataset focused on fine-grained interactions as a complement to existing text-to-image benchmarks. While most of the previous approaches failed to generate accurate details, we proposed DetailScribe, a generate-then-refine framework that leverages hierarchical critiques from vision-language models to iteratively refine text-to-image generations. By breaking down prompts into structured hierarchies and utilizing VLM feedback to guide a diffusion-based refinement process, our approach effectively improves text-to-image generation tasks with fine-grained details of entity interactions. We evaluate different algorithms for text-to-image generation and refinement on our interaction-rich dataset InterActing, and demonstrate that our model DetailScribe achieves superior semantic accuracy and visual coherence.
Supplementary Material
The supplementary material is organized as follows. In Appendix A, we describe the text prompts for the vision language model (GPT-4o) during the concept decomposition stage and the critique-and-refinement generation stage. In Appendix C, we provide quantitative automatic evaluation per scenario, and examples for the comparison of different metrics. In Appendix D, we provide more qualitative examples generated by DetailScribe, more qualitative examples for the ablation studies, and additional discussions on the failure modes of the system. Finally, in Appendix E, we describe the full dataset of our text prompts.
@@ A.1 Prompt: Concept Decomposition
The following is the input prompt to GPT-4 model for the concept decomposition task. Throughout the entire dataset, we consistently use the single example of “cooking” as the single-shot example.
@@
1. Topic: polar-bear-cut-cake Prompt: An anime of a polar bear carefully cutting a berry cake. Components: (concept: polar-bear-cut-cake) = (paw hold knife) + (knife cut through cake) + (cake rest on plate) 2. Topic: Inside a Circle Prompt: A circle of sunflowers with a single, vibrant red rose in the very center, surrounded by the larger yellow blooms. Components: (concept: circle) = (sunflowers form circle) + (red rose placed in center) + (yellow blooms surround rose)”, ”description”: ”A circle of sunflowers with a single, vibrant red rose in the very center, surrounded by the larger yellow blooms.
@@ A.2 Prompt: Critique and Prompt Refinement
Appendix B InterActing Prompting Implementation Details
@@ B.1.3 Prompt Completion
Come up with a description of an animal [content], the description should be similar as the following example and uncommon to be observed. Do not use passive voice. Double check the description to focus on major relation, which is [content]. Write down your answer in this format: {”topic”: [content], ”prompt”: description} For example: interaction: ”taking photos” Entities: squirrel Description: A squirrel taking photos with a camera. Then, the output should be: {”topic”: ”taking-photos”, ”prompt”: ”A squirrel taking photos with a camera.”}
B.2 Compositional Spatial Relationships
@@ B.2.1 Topic Generation
B.2.2 Prompt Completion
Come up with a description of a scene which is a novel combination of [content], the description should be similar as the following example and uncommon to be observed. Do not use passive voice. Double check the description to focus on major relation, which is [content]. Write down your answer in this format: {”topic”: [content], ”prompt”: description} For example: Concept: ”tic-tac-toe” Description: A tic-tac-toe composed by tomato and cucumber as the players symbols. Then, the output should be: {”topic”: ”tic-tac-toe”, ”prompt”: ”A tic-tac-toe composed by tomato and cucumber as the players symbols.”}
B.3 Multi-subject Interactions
B.3.1 Topic Generation

### v2

@@ 5 Experiments
We compare DetailScribe with state-of-the-art text-to-image generation models on the InterActing dataset. Additionally, we carry out ablation studies on the incorporation of hierarchical concept decomposition and the number of re-denoising steps.
Functional Relation Multi-subject Interaction Compositional Relation
Human Qwen2.5 ImReward CLIPS. B-VQA Human Qwen2.5 ImReward CLIPS. B-VQA Human Qwen2.5 ImReward CLIPS. B-VQA
SD 3.360 4.480 1.657 0.971 0.366 3.225 3.800 1.171 0.894 0.306 3.583 4.867 1.415 0.898 0.379
+ GPT Rewrite 3.770 4.440 1.546 0.921 0.245 3.275 3.900 1.101 0.890 0.292 3.167 4.467 1.283 0.867 0.325
+ GPT Refine 3.450 4.600 1.524 0.951 0.290 3.175 3.600 1.022 0.858 0.292 3.667 4.800 1.390 0.889 0.389
+ Inf Scale 3.270 4.560 1.718 0.985 0.434 3.375 4.400 1.149 0.903 0.302 3.650 4.800 1.538 0.912 0.423
DALL·E 3 3.940 4.680 1.535 0.880 0.226 3.775 4.700 1.111 0.813 0.286 3.433 4.867 1.382 0.838 0.367
DetailScribe 4.280 4.880 1.761 0.998 0.449 3.800 4.600 1.326 0.907 0.343 4.283 5.000 1.545 0.923 0.485
Table 2: Average human/MLLM likert scale (1 - 5) and pre-trained metrics on three scenarios of sampled InterActing dataset. We report the human Likert scale (Human Evaluation), MLLM evaluation score (Qwen2.5-VL-32B), as well as ImageReward (ImReward), CLIPScore (CLIPS.) and BLIP-VQA (B-VQA) score. DetailScribe receives the highest scores according to human preference in all scenarios.
@@ 5.1 Baselines and Implementation Details
We compare our model with the following models:
  • Stable Diffusion. We generate the image conditioned on the prompts in InterActing using the Stable Diffusion SD3.5-large model (SD). We used the SD3.5-large model Esser et al. (2024) for all the baselines that are based on a pre-trained T2I generative model.
  • Refinement-Augmented Generation We also include two common strategies for refinement-augmented generation (SD + GPT Rewrite and SD + GPT Refine). For SD + GPT Rewrite, we first use GPT-4o to generate a detailed text prompt given the initial prompt from InterActing, akin to the concept decomposition used in DetailScribe. This improved text prompt is then used to re-generate images with the same pre-trained T2I model. For SD + GPT Refine, we follow the strategy in Saroj and Patel (2025). Specifically, GPT-4o was adopted to provide a refined prompt based on an initially generated image and the refinement request prompt which is used in Saroj and Patel (2025). We then re-generated the image with the refined prompt.
  • Inference scaling  (Ma et al., 2025). We implemented a toy version of noise searching in  Ma et al. (2025) (SD + Inf Scale). Specifically, we sampled two different noises and adopted CLIPScore (Hessel et al., 2022) as its verifier.
  • DALL·E Ramesh et al. (2022); Bubeck et al. (2023) internally integrates LLMs (GPT) to refine prompts with detail model interprets effectively before generating images. We include the DALL·E 3 as a strong baseline to assess DetailScribe’s advancement in interpreting and generating scenes with rich entity interactions.
  • DetailScribe. Our DetailScribe implementation leverages Stable Diffusion 3.5 as the foundational model for both image generation and refinement. To ensure fair comparisons, all approaches involving MLLMs and LLMs use separate, identical prompts with the same GPT-4o model, maintaining consistency across evaluations.
Among all the algorithms, SD + GPT Refine and inference scaling (SD + Inf Scale), as well as DetailScribe, require two times the computation of the base model (SD) for one iteration of refinement (We neglected the critique time of the MLLM done by commercial APIs, which has a runtime of 10% of SD3.5.) SD + GPT Rewrite refines prompt unconditional on previous generation, thus requires the same computation as SD.
@@ 5.2 Result
We evaluate the models on three scenarios from the InterActing dataset and report the results separately. Due to the scalability of high-quality human evaluation, we sampled 50 prompts from InterActing for both human evaluation and automatic evaluation (Table 2), and compared the agreement in between. We adopted Qwen2.5-VL-32B (Bai et al., 2025) as our MLLM evaluator to avoid evaluation bias. Overall, MLLM evaluator achieves the highest agreement at 87.6%, compared to the other metrics: ImageReward (73.6%), CLIPScore (70.4%), and BLIP-VQA (67.6%). We further presented the automatic evaluation on the entire InterActing in Table 3. DetailScribe outperformed all methods based on SD3.5 in all evaluation.
Qwen2.5 ImReward CLIPS. B-VQA
SD 4.228 1.323 0.902 0.336
+ GPT Rewrite 4.116 1.193 0.880 0.268
+ GPT Refine 4.141 1.255 0.880 0.300
+ Inf Scale 4.126 1.354 0.922 0.365
DALL·E 3 4.519 1.222 0.860 0.312
DetailScribe 4.570 1.460 0.923 0.381
Table 3: Automatic evaluation on entire InterActing. DetailScribe outperforms all baselines on all metrics. We include evaluation by scenario in Appendix D.
Fig. 4 shows more examples generated by our model and the baselines. As shown in the figure, DetailScribe is able to generate images with fine details delineating entity interactions. For example, the first row demonstrates the capability of DetailScribe in capturing functional relations. Given the prompt “A cat sails across the sea in a large seashell, holding a mast.”, all baselines fail to capture the relation “holding”, while DetailScribe is able to generate accurate details of a cat holding the mast with fine details.
The 4th row contains a challenging example of a complex scene layout “zig-zag path”. DetailScribe is the only model capable of generating such fine layout patterns. Stable Diffusion can generate an image with zigzag patterns but fails to reveal a path. Both SD with GPT rewritten and refined prompt and DALL·E fail to follow the prompt on the zigzag pattern and only generate a path with leaves.
@@ 5.3 Ablation: Hierarchical Concept Decomposition Improves Error Detection
Figure 5: An illustrative example showing the effectiveness of the explicit concept decomposition module. MLLM first critiques the original generation, and identifies the features needs to be correct (red) and the features non-necessary for further modification (green), and then provides the corrected prompt for re-denoising.
@@
Fig. 5 shows an illustrative example. The top box in the figure shows critiques generated without our decomposed concepts. The MLLM with access to only the initial prompt does not properly attend to the detailed properties. Instead, the critiques focus more on global attributes such as the shape of objects, lighting conditions, or object arrangements. With the concept decomposition step explicitly added, the MLLM can generate better critiques by attending to local details such as “missing a spoon” or action concepts such as “stirring”. With more errors detected and included in the refined prompt, we also see an improvement in the re-denoising image refinement.
Human Qwen2.5 ImReward CLIPS. B-VQA
w/o. Decomp 3.843 4.660 1.586 0.953 0.410
w/. Decomp 4.187 4.760 1.609 0.957 0.438
Table 4: Ablation study on the effectiveness of the concept decomposition module. Explicit concept decomposition significantly improves the generation quality.
@@ 5.4 Ablation: Re-Denoising Step
Figure 6: Ablation on the number of re-denoising steps. The first column shows the result without the refined prompt. Column 2-5: Starting the re-denoising at step $T$, T1T-1, T4T-4, and T6T-6.
We also study the impact of using different numbers of re-denoising steps, as shown in Fig. 6. There are tradeoffs between re-denoising steps and editing level. When $t^{\prime}$ is set to a very large value (e.g., $T$, corresponding to complete noise), although we can correct the identified errors, the entire image is effectively regenerated. This often introduces new errors. By contrast, if $t^{\prime}$ is set too low (i.e., when the image is already nearly clean), the updates from the refined prompt have minimal impact, making it difficult to incorporate necessary corrections. To study the tradeoffs, We generate images with SD3.5, starting from steps $T$, T1T-1, T4T-4, and T6T-6, respectively. We ablate the choice of $t^{\prime}$ in detail in the experiment section. Introducing the refinement prompt at a later de-noising step results in images that are more similar to the original ones, as the diffusion model has fewer steps to refine the generated results. Overall, we find that salient and global attributes of objects, such as the shapes and colors of large objects, are less likely to be modified if the re-denoising occurs at a late stage. However, it is still possible to make local changes that do not interact with large regions of the image, such as adding small objects. In general, for Stable Diffusion models, we found that setting $t^{\prime}$ close to T2T-2 gives the best performance.
We also evaluate the performance of generating images from pure noise using the refined prompt. Given that the refined prompts contain more details, we empirically observe that diffusion models are more prone to missing concepts or concept leakage, such as missing entities (e.g. ignoring “Snow Bunny” given the prompt “A Rabbit Sculpting Snow Bunny”). Thus, our refinement-augmented generation procedure can also be interpreted as a coarse-to-fine generation process, making it easier for the model to generate images with coherent global structures and detailed local attributes simultaneously. Furthermore, while we find that a one-round refinement is sufficient, future work may explore extending the framework to multi-round refinements.
@@ 6 Conclusion
In this paper, we introduce InterActing, a comprehensive dataset focused on fine-grained interactions as a complement to existing text-to-image benchmarks. While most of the previous approaches failed to generate accurate details, we proposed DetailScribe, a generate-then-refine framework that leverages hierarchical critiques from MLLMs to iteratively refine text-to-image generations. We evaluate different algorithms for T2I generation and refinement on our interaction-rich dataset InterActing, and demonstrate that our model DetailScribe achieves superior semantic accuracy and visual coherence.
Limitation. To ensure the effectiveness of refinement, we prevent introducing significant changes to the image content by avoiding re-denoising from complete noise. A current limitation of the DetailScribe framework, therefore, is its reliance on that the image generated without prompt refinement has a correct global scene structure. For example, if the generated image based on the user input completely misses one of the main subjects. In such cases, even if the MLLM detects errors in the generated image, the re-denoising process may not fix them. Future work may explore seed search Ma et al. (2025) based on a similar critique strategy.
Acknowledgements
We thank Xiaolin Fang, Lei Xu and Lijie Fan for helpful discussion and proofreading. We thank Bowen Pan, Jiejun Jin, Wendi Gao and Yifan Le for volunteering high-quality human evaluation.
We gratefully acknowledge support from NSF grant 2214177; from AFOSR grant FA9550-22-1-0249; from ONR MURI grant N00014-22-1-2740; and from ARO grant W911NF-23-1-0034; from MIT Quest for Intelligence; from the Boston Dynamics AI Institute; from ONR Science of AI; and from Simons Center for the Social Brain. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of our sponsors.
Appendix
The appendix is organized as follows. In Appendix A, we describe the text prompts for the vision language model (GPT-4o) during the concept decomposition stage and the critique-and-refinement generation stage. In Appendix B, we show the text prompts for interaction topic sampling and completion of InterActing. In Appendix C, we provide details of evaluation guidance, the quantitative automatic evaluation per scenario, and examples for the comparison of different metrics. In Appendix D, we provide more qualitative examples generated by DetailScribe and additional discussions on the failure modes of the system. Finally, in Appendix E, we describe the full dataset of our text prompts.
@@ A.1 Prompt: Concept Decomposition
The following is the input prompt to GPT-4o model for the concept decomposition task. Throughout the entire dataset, we consistently use the single example of “cooking” as the single-shot example.
@@
1. Topic: polar-bear-cut-cake Prompt: An anime of a polar bear carefully cutting a berry cake. Components: (concept: polar-bear-cut-cake) = (paw hold knife) + (knife cut through cake) + (cake rest on plate) 2. Topic: Inside a Circle Prompt: A circle of sunflowers with a single, vibrant red rose in the very center, surrounded by the larger yellow blooms. Components: (concept: circle) = (sunflowers form circle) + (red rose placed in center) + (yellow blooms surround rose)", "description": "A circle of sunflowers with a single, vibrant red rose in the very center, surrounded by the larger yellow blooms.
@@ A.2 Prompt: Critique and Prompt Refinement
Appendix B InterActing Generation Details
@@ B.1.3 Prompt Completion
Come up with a description of an animal [content], the description should be similar as the following example and uncommon to be observed. Do not use passive voice. Double check the description to focus on major relation, which is [content]. Write down your answer in this format: {"topic": [content], "prompt": description} For example: interaction: "taking photos" Entities: squirrel Description: A squirrel taking photos with a camera. Then, the output should be: {"topic": "taking-photos", "prompt": "A squirrel taking photos with a camera."}
B.2 Multi-subject Interactions
@@ B.2.1 Topic Generation
B.2.2 Prompt Completion
Come up with a description of two animals doing [content], the description should be similar as the following example and uncommon to be observed. Do not use passive voice. Double check the description to focus on major relation, which is [content]. Write down your answer in this format: {"topic": [content], "prompt": description} The description must contains the exact [content] word. For example: Concept: "High-Fiving" Description: A dolphin and a seal leap from the water, high-fiving with their flippers. Then, the output should be: {"topic": "High-Fiving", "prompt": "A dolphin and a seal leap from the water, high-fiving with their flippers."}
B.3 Compositional Spatial Relationships
B.3.1 Topic Generation